OpenStudy (anonymous):

Using complete sentences, explain how you would graph the equation y + 4x = 8 (3 pts.). Explain how you would make a table of values. List the values you would use to graph the equation

OpenStudy (texaschic101):

y + 4x = 8 to find the x intercept, sub in 0 for y. 0 + 4x = 8 4x = 8 x = 2 ......the x intercept is (2,0) to find the y intercept, sub in 0 for x y + 4(0) = 8 y = 8 ..... the y intercept is (0,8) plot your intercepts (2,0) and (0,8) you can then draw a line connecting your intercepts. But if you need to use the slope, put your equation in y = mx + b form y + 4x = 8 y = -4x + 8 slope is -4. start at (0,8) and go down 4 spaces, and to the right 1 space, then down 4, then to the right 1, and you should cross the x axis at (2,0) ========================= as far as making a table... y + 4x = 8 when x = 1 y + 4(1) = 1 y + 4 = 1 y = -4 + 1 y = -3 when x = 2 y + 4(2) = 1 y + 8 = 1 y = -8 + 1 y = -7 when x = 3 y + 4(3) = 1 y + 12 = 1 y = -12 + 1 y = -11 and so on....depending on how many values you want in your table.